I am trying to help my daughter with her home work and we are having issues with this problem.
The length of a rectangle is 150% of its width, w. Use the percent equation to find the length of the rectangle in terms o fw. Then write a formual for the area of the rectangle using only the variable w.

Answer by richwmiller(17219)

L*W=Area
L=1.5W
1.5W*W=A
1.5W^2=A